Find the height of a cuboid whose base area is 180 cm^2 and volume is 900 cm^3?

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Given, Base area of cuboid = 180 cm2and Volume of cuboid = 900 cm3
We know that, Volume of cuboid = lbh
900 = 180×h (substituting the values)
h= 900/180 = 5
Hence the height of cuboid is 5 cm.

Thanks